Factor the following polynomial out completely.
-3s^2-10s+8

Answer by checkley75(3666)

-3s^2-10s+8=0
(-3s+2)(s+4)=0
-3s+2=0
-3s=-2
s=-2/-3
s=2/3 answer.
s=4=0
s=-4 answer.